OpenStudy (anonymous):

Two stores sell the same computer for the same original price. Store A advertises that the computer is on sale for 25% off the original price. Store B advertises that it is reducing the computer’s price by $180. When Brittany compares the sale prices of the computer in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the computer’s original price. A. 0.25p = p – 180 B. 0.75p = p + 180 C. 0.25(p – 180) = p D. 0.75p = p − 180

OpenStudy (anonymous):

I think its D but is i'm not right explane how you got the answer please thank you... OH AND I GIVE MEDALS!

OpenStudy (anonymous):

@Rockiee

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@samson245

OpenStudy (anonymous):

@bazel_e

OpenStudy (anonymous):

c i think not sure @3spooky5me might be able to help

OpenStudy (anonymous):

youre right

OpenStudy (anonymous):

d is correct

OpenStudy (anonymous):

can you explane how you guys got the answer xD

OpenStudy (anonymous):

d is correct oh

OpenStudy (anonymous):

basically you simply find out what the percentage reduction is as a price, then see which answer contains the proper % of price drop to equal the second stores direct $ advertisement.

OpenStudy (anonymous):

nvm then lol i didn't see you say that it was D lol

OpenStudy (anonymous):

p - 25% =p - 180 p - 0.25p = p -180 0.75p=p-180

OpenStudy (anonymous):

oh aha

OpenStudy (anonymous):

sorry